1. A black hole with a mass of 1.1 x 10^31 kg pulls a planet towards itself with a force of gravity of 2.8 x 10^21  If the distance between their centres of mass is 3.5 x 10^11 what is the mass of the planet?
1
Expert's answer
2021-01-12T12:11:57-0500
"F=G\frac{Mm}{r^2}\\2.8 \cdot 10^{21}=(6.67\cdot 10^{-11})\frac{(1.1\cdot 10^{31})m}{(3.5\cdot 10^{11})^2}\\m=4.7\cdot 10^{23}kg"